Two years ago, achillies was half as old as Ajax will be three years from now. Achilles is now 17. How old is Ajax now
Answer by stanbon(75887)

Two years ago, achillies was half as old as Ajax will be three years from now. Achilles is now 17. How old is Ajax now
---
Equations:
Achil = 17
---
(Achil-2) = (1/2)(Aj+3)
---
Substitute for Achil and solve for "Aj":
17 = (1/2)(Aj + 3)
34 = Aj + 3
Ajax = 31 years old now
==============
Cheers,
Stan H.
===============